OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). )y − 4 = –1\2 (x + 6) y – 4 = 2(x + 6) y + 6 = –1\2 (x − 4) y + 6 = 2(x – 4)

OpenStudy (anonymous):

use the point-slope equation\[y=m(x-x_1)+y_1\]and the slope formula\[m=\frac{y_2-y_1}{x_2-x_1}\]like this:

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{4-0}{-6-2}=-\frac{1}{2}\]

OpenStudy (anonymous):

\[y=m(x-x_1)+y_1\]\[y=-\frac{1}{2}(x-2)+0\]\[y=-\frac{1}{2}x+1\]

OpenStudy (radar):

That would be the first one. y=-1/2(x+6) +4 y=(-1/2) (x) -3 +4 y=(-1/2) x + 1